Assume that a, b, c, d are positive integer such that a to the power 5=b  t the power 4
And c to the power 3=d to the power 2 and c-a=19: .  Then find out d-b

Grade:10

1 Answers

Arun
5 years ago
Dear Trishna
 
a^5 = b^4
 
now assume a = alpha^4
 
hence b = alpha^5
 
similrily
 
c^3 = d^2
 
assume
 
c = beta^2
 
then
d = beta^3
 
Hope now you can solve it